OpenStudy (anonymous):

The length of a rectangular piece of land is 35ft more than four times its width. The perimeter is 770ft. Find its dimensions.

OpenStudy (campbell_st):

well let the width be w then the length is l = w + 35 and the perimeter is 770 = 2 l + 2 w substitute the value for l 770 = 2(w + 35) + 2w simplify and solve for w when you have w use l = w+ 35 to find the length

OpenStudy (anonymous):

Thanks a lot :)